# To reset or not to reset in a finite domain: that is the question

### v1

@@ Abstract
We investigate the search of a target with a given spatial distribution in a finite one-dimensional domain. The searcher follows Brownian dynamics and is always reset to its initial position when reaching the boundaries of the domain (boundary resetting). In addition, the searcher may be reset to its initial position from any internal point of the domain (bulk resetting). Specifically, we look for the optimal strategy for bulk resetting, i.e., the spatially dependent bulk resetting rate that minimizes the average search time. The best search strategy exhibits a second-order transition from vanishing to non-vanishing bulk resetting when varying the target distribution. The obtained mathematical criteria are further analyzed for a monoparametric family of distributions, to shed light on the properties that control the optimal strategy for bulk resetting. Our work paves new research lines in the study of search processes, emphasizing the relevance of the target distribution for the optimal search strategy, and identifies a successful framework to address these questions.
@@
Nature is full of instances of first-passage processes [1, 2]. Search problems can be formulated in terms of them, focusing on analyzing specific quantities related to the first-passage time—like the mean first-passage time (MFPT)—as key indicators of search performance. These problems are especially relevant from an applied perspective and have a significant impact across a wide range of disciplines, from biophysics—at various scales, from single molecules [3] to animal foraging [4, 5]—to computing methods [6], rescue mission strategies [7], or reliability theory [8, 9], to name just a few.
@@
The typical goal in a search problem is the optimization of the search strategy, e.g., finding out the strategy that minimizes the MFPT [10]. Intermittent strategies that combine reactive random search phases with faster, non-reactive relocation phases, have proven to be highly effective to this end [11]. In its simplest introduction, stochastic resetting constitutes a specific instance of an intermittent strategy, with instantaneous relocations [12, 13]. In the last years, optimization of stochastic resetting has been a prolific line of research [14, 13, 15, 16, 17, 18, 9, 19, 20, 21, 22, 23, 24]. The original simple proposal for stochastic resetting has been sophisticated by implementing more realistic reset mechanisms, e.g., using confining potentials [25, 18, 26, 27, 20, 28] or introducing some cost for the resets [29, 30, 9, 31, 32, 33].
Remarkably, most of optimization problems within search research have been posed for a target with a well-defined position, with very few exceptions [10, 13, 23]. Yet, optimizing the search strategy for a fixed target position is basically pointless for a practical search problem: the optimal strategy thus found depends on the target position, of which an actual searcher is uncertain. Therefore, the role of the degree of uncertainty about the target location—codified in its spatial distribution—is a significant issue that demands to be investigated to develop optimal strategies in more realistic contexts. Moreover, it must be remarked that stochastic resetting has been predominantly studied in infinite domains, where resets are always beneficial because they “kill” the trajectories drifting away from the target. For finite domains [34, 35, 36, 9, 37], resetting has been observed to be detrimental for some specific situations.
@@
In this Letter, we address the gaps described above by analyzing the impact on the optimal resetting strategy of the target’s spatial distribution in a finite domain. To this end, we employ the paradigmatic model of a resetting Brownian searcher moving in a finite one-dimensional box. In this context, we will look for the best resetting strategy, allowing for it to be inhomogeneous, i.e., the resetting rate depends in general on the position of the searcher. Space-dependent resetting has been considered in systems with a localized target [14], and the stationary state obtained by using a quantum mechanical mapping [38].
For the considered one-dimensional search, a bounded distribution naturally leads to the concept of resetting boundaries: if the searcher reaches either wall of the box without finding the target, the optimal strategy is clearly to restart the search from its initial position. Finally, each target distribution determines an optimal—potentially inhomogeneous—bulk resetting rate. Here, we show how this optimal rate can be derived from the target distribution, and we provide clear criteria for when bulk resetting can be avoided.
@@
Model.- We consider a Brownian searcher that freely diffuses in a one-dimensional box. Let $X(t)$ be its position at time $t$, which is looking for the target located at $X_{T}$. From a mathematical point of view, $X$ is a stochastic process and $X_{T}$ is a random variable with a certain distribution $P$. The distribution $P$ quantifies the degree of uncertainty about the location of the target; flat and single delta-peaked distributions, respectively, represent extreme cases of full uncertainty and full certainty. The target remains motionless in any trajectory, which has been referred to as a quenched-disorder target [23] due to its similarity to disordered systems  [^1] As opposed to an annealed-disorder target, whose position would evolve in time..
@@
The searcher starts from the center of the box, $x=0$. When it reaches the target at $x=x_{T}$, the dynamics terminates. Until then, its dynamics—illustrated in Fig. 1—is as follows: (i) When hitting the boundaries at $x=\pm\ell$, the position is instantaneously reset to the center; (ii) the probability per unit time $r$ of a reset event in the bulk depends solely on space, i.e., $r\equiv r(x)$, and the reset event instantaneously restarts the search from the center; (iii) between reset events, the searcher freely diffuses with a constant diffusion coefficient $D$.
@@
On the one hand, our resetting boundaries clearly expedite—with respect to reflecting ones—the search in a one-dimensional domain: reaching one wall without success means that the target lies on the other half of the box. On the other hand, bulk resets at a certain position $x$ are beneficial (detrimental) if $x\,x_{T}$ is negative (positive). Therefore, two natural questions arise: When is bulk resetting advantageous for finding the target, given its distribution $P(x_{T})$? What is the optimal bulk resetting strategy $r(x)$ for a given $P(x_{T})$? The main goal of this Letter is to answer these two key questions.
For search optimization, we start by studying the MFPT $\tau(x_{0}|x_{T})$ from an arbitrary initial position $x_{0}$ to the target at a fixed location $x_{T}$. Specifically, we focus on the MFPT $\tau_{0}(x_{T})\equiv\tau(0|x_{T})$ from the center of the box to the target. Then, we consider its average with respect to the target distribution
[…]
where we have introduced the notation $\langle\cdot\rangle_{T}$ for averages with respect to $P(x_{T})$. In this Letter, the average time $\overline{\tau}[r]$ is the main figure of merit for quantifying the success of the search. Note that $\overline{\tau}[r]$ is a functional of the bulk resetting strategy $r(x)$ through $\tau_{0}$—the functional dependence of the latter on $r(x)$ has not been explicitly included in the notation to avoid cluttering the formulas.
The MFPT at fixed $x_{T}$ can be obtained by solving the backward Fokker-Planck equation
@@
where we have switched to dimensionless variables, respectively taking $\ell$ and $\ell^{2}/D$ as units for space and time, as detailed in Sec. I of [40]. The boundary conditions (2b) implement resetting boundaries at the walls and the absorbing point at the target position, respectively. Since we are interested in $\tau_{0}(x_{T})$, it suffices to solve Eq. (2) for $x_{0}$ in the interval between one of the walls and the target position that includes $x=0$. While the analytical solution for any bulk resetting rate $r(x)$ is out of reach, it can still be obtained for specific choices of $r(x)$.
Homogeneous resetting.- Here, we consider the average MFPT defined in Eq. (1) for a homogeneous bulk resetting rate $r$. The average MFPT is then no longer a functional, but a standard function of $r$—we thus write $\overline{\tau}(r)$. We will optimize $\overline{\tau}(r)$ for a given $P$, i.e., we look for the best homogeneous resetting rate $r^{*}$ that minimizes $\overline{\tau}$. As already stated above, the dependence on $r$ enters through the solution for $\tau_{0}$ from Eq. (2), which can be solved for constant $r$—see Sec. II of [40]:
@@
To obtain $\overline{\tau}(r)$, it suffices to compute the average of Eq. (3) with respect to $P(x_{T})$. We note that $\tau_{0}(x_{T})$ is a convex function that diverges for $r\to\infty$. Since $r\geq 0$, the behavior of $\overline{\tau}(r)$ close to $r=0$ depends on the sign of $d\overline{\tau}/dr|_{r=0}$. For positive derivative, $\overline{\tau}(r)$ has a local minimum at $r=0$, so the best search strategy implies no resetting in the bulk, $r^{*}=0$. Instead, for negative derivative, $\overline{\tau}(r)$ has a local maximum at $r=0$ and there appears an optimum non-zero resetting rate, $r^{*}\neq 0$.
@@
Specifically, $m>0$ ($m<0$) entails that the minimum of $\overline{\tau}$ is attained at $r^{*}=0$ ($r^{*}>0$). Using Eq. (3), we get
$M(x_{T})=\frac{1}{24}\left[x_{T}^{4}-\absolutevalue{x_{T}}\left(1-2x_{T}^{2}\right)\right].$ (5)
As shown in the top panel of Fig. 2, $M(x_{T})<0$ for $|x|<x_{T}^{c}\equiv(\sqrt{5}-1)/2$, whereas $M(x_{T})>0$ for $x_{T}^{c}<|x|<1$. Therefore, “small”, close to the center, values of $x_{T}$ contribute to destabilize the no-resetting strategy in the bulk, whereas “large”, close to the walls, values of $x_{T}$ contribute to stabilize it.
Figure 2: Derivative controlling the stability of the no-resetting strategy. Positive (negative) values, represented as warm (cold) colors, thereof contribute to zero (nonzero) bulk resetting as the best search strategy. Top: Standard derivative $M(x_{T})$ defined in Eq. (5) for the homogeneous case. The change of stability is reached at $x_{T}^{c}$ (black diamond). Bottom: Symmetrized functional derivative $M_{s}(x_{T},x)$ defined in Eq. (To reset or not to reset in a finite domain: that is the question) for the inhomogeneous case. White color indicates saturated values. The contour line at which $M_{s}=0$ is marked (dashed line), as well as the value $x_{T}^{(0)}$ (black diamond) below which $M_{s}(x_{T},x)<0$, $\forall x$.
It is worth noting some general results. First, for a flat distribution, $m>0$: in a finite domain, homogeneous resetting does not give any advantage for a homogeneously distributed target. Second, $m<0$ for any target distribution with finite support in the subinterval $0\leq x_{T}<x_{T}^{c}$, because $M(x_{T})<0$ where $P(x_{T})\neq 0$: therein, homogeneous resetting in the bulk is advantageous and $r^{*}\neq 0$. Conversely, $m>0$ for any target distribution with finite support in the subinterval $x_{T}^{c}<x_{T}<1$, because $M(x_{T})>0$ where $P(x_{T})\neq 0$: therein, homogeneous resetting in the bulk is detrimental and $r^{*}=0$.
To further illustrate the discussion above, we now introduce a specific monoparametric family of symmetric target distributions. In particular, we choose a variant of the beta distribution,
@@
defined for $-1<x_{T}<1$, with $\beta>0$. This $\beta$-family interpolates between a single delta peak at the center of the box for $\beta\to+\infty$ and two equally weighted delta peaks at the boundaries for $\beta\to 0$, passing through the flat distribution for $\beta=1$. Introducing Eqs. (3) and (6) into Eq. (1), it is possible to analytically obtain the average MFPT, as detailed in Sec. II of [40]. Yet, for our discussion, it suffices to obtain the linear coefficient $m$ in Eq. (4), which is now a function of $\beta$:
@@
We note that $m(\beta)$ only vanishes at $\beta=\beta_{c}\simeq 1.71$. For $\beta<\beta_{c}$, we have $m>0$ and $\overline{\tau}(r)$ is minimum at $r=0$. Instead, for $\beta>\beta_{c}$, we have $m<0$ and $\overline{\tau}(r)$ has a local maximum at $r=0$. Therefore, the global minimum of $\overline{\tau}(r)$ is attained at a certain value $r^{*}>0$. This behavior is reminiscent of a continuous phase transition, with the optimal homogeneous resetting rate $r^{*}$ playing the role of the order parameter: $r^{*}=0$ for $\beta<\beta_{c}$, $r^{*}\neq 0$ for $\beta>\beta_{c}$, with $r^{*}=0$ for $\beta=\beta_{c}$.
For the homogeneous case, we have proven that nonresetting in the bulk is the optimal strategy for certain target distributions, specifically those for which the region with positive $M(x_{T})$ (close to the walls) is more important, when weighted with the target distribution $P(x_{T})$, than the region with negative $M(x_{T})$ (close to the center). For the monoparametric family (6), this corresponds to $\beta<\beta_{c}$: loosely speaking, the weight of targets close to the center is insufficient to destabilize the strategy with $r=0$ in the bulk. Indeed, this is physically sensible: since the resetting boundaries already facilitate the search, extra resetting in the bulk is beneficial only if the weight of targets close to the center is large enough.
Optimization of heterogeneous resetting.- We now move to the general case with heterogeneous resetting $r(x)$. Similarly to the homogeneous case, whether or not bulk resetting is advantageous is controlled by the derivative of the average MFPT with respect to the resetting rate, evaluated at $r(x)\equiv 0$. However, now this quantity is a functional derivative, specifically

### v2

# How target distributions shape optimal stochastic resetting

@@ Abstract
We investigate the search of a target with a given spatial distribution in a finite one-dimensional domain. The searcher follows Brownian dynamics and is always reset to its initial position when reaching the boundaries of the domain (boundary resetting). In addition, the searcher may be reset to its initial position from any internal point of the domain (bulk resetting). Specifically, we look for the optimal strategy for bulk resetting, i.e., the spatially dependent bulk resetting rate that minimizes the average search time. The best search strategy exhibits a second-order transition from vanishing to nonvanishing bulk resetting when varying the target distribution. The obtained mathematical criteria are further analyzed for different monoparametric families of distributions, which sheds light on the properties that control the optimal strategy for bulk resetting. Our work paves new research lines in the study of search processes, emphasizing the relevance of the target distribution for optimal search strategies, and identifies a successful framework to address these questions.
I Introduction
@@
Nature is full of instances of first-passage processes [1, 2], ranging from the first encounter of two reacting molecules to the first flower visited by a pollinator. Search processes can be formulated in terms of first-passage problems, with the focus on analyzing specific quantities related to the first-passage time—like the mean first-passage time (MFPT)—as key indicators of search performance. These problems are especially relevant from an applied perspective and have a significant impact across a wide range of disciplines, from biophysics—at various scales, from single molecules [3] to animal foraging [4, 5]—to computing methods [6], rescue mission strategies [7], or reliability theory [8, 9], to name just a few.
@@
The typical goal in this context is the optimization of the search strategy, e.g., finding out the strategy that minimizes the MFPT [10]. Intermittent strategies that combine reactive random search phases with faster, nonreactive relocation phases, have proven to be highly effective to this end [11]. In its simplest implementation, stochastic resetting constitutes a specific instance of such intermittent strategies, with instantaneous relocations [12, 13].
In the last years, optimization of stochastic resetting has been a prolific line of research [14, 13, 15, 16, 17, 18, 9, 19, 20, 21, 22, 23, 24]. The original simple proposal for stochastic resetting has been sophisticated by implementing more realistic reset mechanisms, e.g., using confining potentials [25, 18, 26, 27, 20, 28] or introducing some cost for the resets [29, 30, 9, 31, 32, 33].
Remarkably, most optimization problems in search processes have been posed for a target with a well-defined position, with very few exceptions [10, 13, 23, 34]. In such setups, all the results, including optimal strategies, depend on the target position. Nevertheless, the usual situation in a realistic search problem is the searcher being unaware of the target position, for instance, a gatherer looking for food or a rescue mission searching for a lost victim. Therefore, the role of the degree of uncertainty about the target location—codified in its spatial distribution—is a significant issue, worth of being investigated to develop optimal strategies in realistic contexts. This has recently been explored by optimizing the distribution of times between resetting events [34].
Moreover, it must be remarked that stochastic resetting has predominantly been studied in infinite domains, where resets are always beneficial because they impede the trajectories drifting away from the target. For finite domains, the literature is not so vast [35, 36, 37, 9, 38]. In different search setups, where details such as geometry or the number of targets differ, stochastic resetting has been observed to be detrimental in certain situations [35, 36, 37]. In those cases, an optimal nonvanishing resetting rate was found only when the resetting location is close enough to the target position. Regimes where restart—at constant rate—is convenient to accelerate the search fulfill the celebrated CV criterion  [39, 25].
When searches occur in a one-dimensional bounded domain, the advantage of resetting boundaries naturally arises: If the searcher reaches either wall of the box without finding the target, the optimal strategy is clearly to restart the search from its initial position. A typical illustration of stochastic resetting for search processes is how coming back to the starting point may improve the search of your keys at home. When introducing resetting boundaries, we argue that resetting is surely beneficial if the searcher has reached the attic of their house with no success. An analogous idea has been very recently introduced in search processes with multiple searchers [40].
@@
In this article, we address the gaps described above by analyzing how the target’s spatial distribution in a finite domain impacts the optimal resetting strategy. To this end, we employ the paradigmatic model of a resetting Brownian searcher moving in a finite one-dimensional box. Therein, we will look for the best resetting strategy, allowing for it to be heterogeneous: The resetting rate depends in general on the position of the searcher. Our optimization is then different from the recently related work by Evans and Ray [34], where the optimization is carried out over the distribution of times between reset events—keeping the resetting homogeneous. Space-dependent resetting has been previously considered in systems with a localized target [14], and the stationary state has been obtained by using a quantum mechanical mapping [41]. On the one hand, we put our analytical focus on finding whether or not resetting expedites the search process depending on the target distribution. In this regard, we provide neat criteria for determining when bulk resetting is beneficial or detrimental. On the other hand, we illustrate how optimal resetting profiles can be determined for different specific setups.
The rest of the article is organized as follows. In Sec. II, the basics of the resetting Brownian searcher is introduced. We optimize the MFPT over homogeneous resetting rate in Sec. III. In Sec. IV, the resetting strategy is considered heterogeneous. Our focus is to determine whether the absence of bulk resetting constitutes a (local) minimum for the MFPT. Section V is devoted to illustrate specific heterogeneous optimal resetting strategies. We put forward the main conclusions of our study as well as some interesting perspectives in Sec. VI. Finally, some technical derivations and/or discussions can be found in the Appendices.
II Model
We consider a Brownian searcher that freely diffuses in a one-dimensional box.
@@ II Model
Let $X(t)$ be its position at time $t$, and the searcher is looking for the target located at $X_{T}$. From a mathematical point of view, $X$ is a stochastic process and $X_{T}$ is a random variable with a certain distribution $P$. The distribution $P$ quantifies the degree of uncertainty about the location of the target; flat and single delta-peaked distributions, respectively, represent extreme cases of full uncertainty and full certainty. The target remains motionless in any trajectory, which has been referred to as a quenched-disorder target [23] due to the formal similarity with systems with quenched disorder  [^1] As opposed to an annealed-disorder target, whose position would evolve in time..
The essential difference between a distributed target and distributed resetting positions—the latter is not considered in our analysis but has been previously studied [14, 13, 43, 44, 45]—is remarkable. In a single trajectory of the stochastic process, there is only one target drawn from its distribution in the former case, whereas there are multiple resetting locations, drawn from its corresponding distribution, in the latter case. We focus on the problem with a distributed target, motivated by its central role in search processes.
In our model dynamics, the searcher starts from the center of the box, $x=0$, for every trajectory of the stochastic process. When the searcher reaches the target at $x=x_{T}$, the trajectory terminates. Until then, dynamical rules—illustrated in Fig. 1—are as follows:
  1. 1.
    When hitting the boundaries at $x=\pm\ell$, the position is instantaneously reset to the center;
  2. 2.
@@
    the probability per unit time $r$ of a reset event in the bulk depends solely on space, i.e., $r\equiv r(x)$, and the reset event instantaneously restarts the search from the center;
  3. 3.
    between reset events, the searcher freely diffuses with a constant diffusion coefficient $D$.
Hence, the equation governing the forward evolution of the probability distribution $p(x,t|x_{0},t_{0})$ of finding the searcher at position $x$ at time $t$, given that it was at position $x_{0}$ at time $t_{0}$ is
$\partial_{t}p(x,t|x_{0},t_{0})=$ $D\left\{\partial_{x}^{2}p(x,t|x_{0},t_{0})\right$.
$\left.\qquad+\delta(x)\left[\partial_{x}p(x,t|x_{0},t_{0})\right]_{x=\ell}^{x=-\ell}\right\}$
$-r(x)p(x,t|x_{0},t_{0})$
$+\delta(x)\int dx^{\prime}r(x^{\prime})p(x^{\prime},t|x_{0},t_{0}),$ (1)
with initial condition
$p(x,t_{0}|x_{0},t_{0})=\delta(x-x_{0}),$ (2)
and boundary conditions
$p(x_{T},t|x_{0},t_{0})=0,\quad p(\pm\ell,t|x_{0},t_{0})=0.$ (3)
On the right hand side of Eq. (1), the term in the first line stands for standard diffusion, the one in the second line stands for the gain of probability at $x=0$ stemming from the resetting at the boundaries. The terms in the third and fourth lines respectively account for the probability loss at any $x$ and gain at $x=0$ after a reset in the bulk.
Additionally, we have the initial condition for the propagator in Eq. (2), the absorbing condition at the target position $x=x_{T}$ and the resetting boundary conditions at $x=\pm\ell$ in Eq. (3). Since the search instance is terminated when the searcher reaches $x=x_{T}$, for any search instance starting from $x_{0}=0$ the only active boundary is the one opposite to the target, namely $\pm\ell\to-\sigma(x_{T})\ell$ in Eq. (3), being $\sigma$ the sign function. The remaining condition would be always automatically satisfied  [^2] Note that in a generic situation, both conditions (for $x=\pm\ell$) apply. More specifically, if we consider $x_{T}\,x_{0}>0$ and $0<|x_{T}|<|x_{0}|$, both conditions are required..
@@
On the one hand, our resetting boundaries clearly expedite—with respect to reflecting ones—the search in a one-dimensional domain: Reaching one wall without success means that the target lies on the other half of the box. This reasonable fact is checked for a specific setup at the beginning of Sec. III On the other hand, bulk resets at a certain position $x$ are beneficial (detrimental) if $x\,x_{T}$ is negative (positive), since resetting brings the particle closer to (farther from) the target. Therefore, two natural questions arise:
  1. 1.
    When is bulk resetting advantageous for finding the target, given its distribution $P(x_{T})$?
  2. 2.
    What is the optimal bulk resetting strategy $r(x)$ for a given $P(x_{T})$?
The main goal of this study is to answer these two key questions.
Specific limits of the target distribution can already be discussed by relying on physical reasoning and relating them to previous results [23]. Specifically, if the target is certainly at a given position, the best heterogeneous resetting strategy is always to reset ($r(x)\to\infty$) for $|x_{T}-x|>|x_{T}|$ and to suppress resetting for $|x_{T}-x|<|x_{T}|$.
For a general search optimization, we start by studying the MFPT $\tau(x_{0}|x_{T})$ from an arbitrary initial position $x_{0}$ to the target at a fixed location $x_{T}$. Specifically, we focus on the MFPT $\tau_{0}(x_{T})\equiv\tau(0|x_{T})$ from the center of the box to the target. Then, we consider its average with respect to the target distribution
$\overline{\tau}[r]\equiv\expectationvalue{\tau_{0}(x_{T})}_{T}=\int dx_{T}\,\tau_{0}(x_{T})P(x_{T}),$ (4)
where we have introduced the notation $\langle\cdot\rangle_{T}$ for averages with respect to $P(x_{T})$. In this study, the average time $\overline{\tau}[r]$ is the main figure of merit for quantifying the success of the search. Note that $\overline{\tau}[r]$ is a functional of the bulk resetting strategy $r(x)$ through $\tau_{0}$—the functional dependence of the latter on $r(x)$ has not been explicitly included in the notation to avoid cluttering the formulas.
The equation governing the first-passage problem is the backward version of Eq. (1). The MFPT at fixed $x_{T}$ is determined by
@@
where we have switched to dimensionless variables, respectively taking $\ell$ and $\ell^{2}/D$ as units for space and time. The boundary conditions (5b) implement resetting boundaries at the walls and the absorbing point at the target position, respectively.
Since we are interested in $\tau_{0}(x_{T})$, it suffices to solve Eq. (5) for $x_{0}$ in the interval between one of the walls and the target position that includes $x_{0}=0$, that is, either $1,x_{T}]$ for $x_{T}>0$ or $x_{T},1]$ for $x_{T}<0$—analogously to the situation for the boundary conditions for the probability distribution, as discussed after Eq. (3).
Equation (5) is simplified with the change of variables
$F(x_{0})=\tau(x_{0}|x_{T})-\tau(0|x_{T}),$ (6)
which removes the source term $\tau(0|x_{T})$,
$-1$ $=\partial_{x_{0}}^{2}F(x_{0})-r(x_{0})F(x_{0})\ ,$ (7a)
$F(0)$ $=0,\quad F(\pm 1)=0\ .$ (7b)
The dependence on $x_{T}$ is not explicitly written from now on to make our notation more concise. Since we are interested in the MFPT from the origin to the target position, once we solve the differential problem for $F$, we need to evaluate the solution at $x_{0}=x_{T}$, i.e., $\tau_{0}(x_{T})=-F(x_{T})$. As for Eq. (3), the boundary conditions (7b) make the system overconstrained; although they are both valid, Eq. (7) will be solved only by activating the boundary condition corresponding to $x_{0}=-\sigma(x_{T})$. While the analytical solution for any bulk resetting rate $r(x)$ is out of reach, it can still be obtained for specific choices of $r(x)$.
III Optimization over homogeneous resetting
We here consider the average MFPT defined in Eq. (4) for a homogeneous bulk resetting rate $r$. Since resetting boundaries have been proved to be beneficial for the search problem, we will always assume their presence—even when bulk resetting vanishes, i.e., $r=0$. The average MFPT is then no longer a functional, but a standard function of rr—we thus write $\overline{\tau}(r)$. We will optimize $\overline{\tau}(r)$ for a given $P$, i.e., we look for the best homogeneous resetting rate $r^{*}$ that minimizes $\overline{\tau}$. As already stated above, the dependence on $r$ enters through the solution for $\tau_{0}$ from Eq. (5), which can be solved for constant $r$,
@@ III Optimization over homogeneous resetting
Had we consider reflecting boundaries instead of our resetting boundaries, the result would have been
$\tau^{\text{(refl)}}_{0}(x_{T};r)=r^{-1}\left[\frac{\cosh[(|x_{T}|+1)\sqrt{r}]}{\cosh[\sqrt{r}]}-1\right]\ ,$ (9)
where we have taken into account that boundary conditions have changed. Specifically, the MFPT still verifies Eq. (7a), but the boundary conditions for $x_{0}=\pm 1$ in Eq. (7b) are substituted with $\left.\partial_{x_{0}}F(x_{0})\right|_{x_{0}=\pm 1}=0$. As physically reasonable, it is straightforward to check that $\tau^{\text{(refl)}}_{0}(x_{T};r)\geq\tau_{0}(x_{T};r)$, $\forall(x_{T};r)$: This constitutes a mathematical proof of the advantage of resetting boundaries over reflecting ones. Henceforth, we focus on the setup with resetting boundaries.
To obtain $\overline{\tau}(r)$, it suffices to compute the average of Eq. (8) with respect to $P(x_{T})$. We note that $\tau_{0}(x_{T};r)$ is a convex function of $r$, $\partial_{r}^{2}\tau_{0}(x_{T};r)>0$, which diverges for $r\to\infty$. Convexity is inherited for the average MFPT $\overline{\tau}(r)$ for arbitrary $P(x_{T})$. Since $r\geq 0$ by definition, the behavior of $\overline{\tau}(r)$ close to $r=0$ depends on the sign of $d\overline{\tau}/dr|_{r=0}$. For positive derivative, the global minimum of $\overline{\tau}(r)$ is located at r=0r=0—due to the already discussed convexity: Therefore, the best search strategy implies no resetting in the bulk, $r^{*}=0$. Instead, for negative derivative, the global minimum of $\overline{\tau}(r)$ is no longer located at $r=0$: There appears an optimum nonzero resetting rate, $r^{*}\neq 0$; recall that $\overline{\tau}(r)$ diverges for $r\to\infty$.
@@
Specifically, $m>0$ ($m<0$) entails that the minimum of $\overline{\tau}$ is attained at $r^{*}=0$ ($r^{*}>0$). The analysis of the slope at $r=0$ has been successfully used in the past for optimizing different resetting configurations [35, 17, 47, 48]. Using Eq. (8), we get
$M(x_{T})=\frac{1}{24}\left[x_{T}^{4}-\mathinner{\!\left\lvert x_{T}\right\rvert}\left(1-2x_{T}^{2}\right)\right].$ (11)
As shown in the top panel of Fig. 2, $M(x_{T})<0$ for $|x_{T}|<x_{T}^{c}\equiv(\sqrt{5}-1)/2$, whereas $M(x_{T})>0$ for $x_{T}^{c}<|x_{T}|<1$. Therefore, “small”, close to the center, values of $x_{T}$ contribute to destabilize the nonresetting strategy in the bulk, whereas “large”, close to the walls, values of $x_{T}$ contribute to stabilize it.
Figure 2: Derivative controlling the stability of the nonresetting strategy for homogeneous resetting, i.e., $M(x_{T})$ defined in Eq. (11). Positive (negative) values, represented as warm (cold) colors, thereof contribute to zero (nonzero) bulk resetting as the best search strategy. The change of stability is reached at $x_{T}^{c}$ (black diamond).
It is worth noting some general results. First, for a flat distribution, $m>0$; in a finite domain, homogeneous resetting does not give any advantage for a homogeneously distributed target. Second, $m<0$ for any target distribution with finite support in the subinterval $0\leq x_{T}<x_{T}^{c}$, because $M(x_{T})<0$ where $P(x_{T})\neq 0$: Therein, homogeneous resetting in the bulk is advantageous and $r^{*}\neq 0$. Conversely, $m>0$ for any target distribution with finite support in the subinterval $x_{T}^{c}<x_{T}<1$, because $M(x_{T})>0$ where $P(x_{T})\neq 0$: Therein, homogeneous resetting in the bulk is detrimental and $r^{*}=0$. The discussion is analogous for $1<x_{T}<0$ since Eq. (11) is an even function of $x_{T}$.
To further illustrate our discussion, we now introduce a specific monoparametric family of symmetric target distributions. In particular, we choose a variant of the beta distribution,
@@
defined for $1<x_{T}<1$, with $\beta>0$. This $\beta$-family interpolates between a single delta peak at the center of the box for $\beta\to+\infty$ and two equally weighted delta peaks at the boundaries for $\beta\to 0$, passing through the flat distribution for $\beta=1$. Since its support is the whole domain [1,1][-1,1], the general observations made above for especially supported distributions do not apply because there is a competition between positive and negative contributions to $M(x_{T})$. Introducing Eqs. (8) and (12) into Eq. (4), it is possible to analytically obtain the average MFPT, as detailed in Appendix A.
Yet, for our discussion, it suffices to obtain the linear coefficient $m$ in Eq. (10), which is now a function of $\beta$:
@@
We note that $m(\beta)$ only vanishes at $\beta=\beta_{c}\simeq 1.71$. For $\beta<\beta_{c}$, we have $m>0$ and the global minimum of $\overline{\tau}(r)$ is found at $r=0$. Instead, $m<0$ for $\beta>\beta_{c}$.
Therefore, the global minimum of $\overline{\tau}(r)$ is attained at a certain value $r^{*}>0$. This behavior is reminiscent of a continuous phase transition, with the optimal homogeneous resetting rate $r^{*}$ playing the role of the order parameter: $r^{*}=0$ for $\beta<\beta_{c}$, $r^{*}\neq 0$ for $\beta>\beta_{c}$, with $r^{*}=0$ for $\beta=\beta_{c}$.
For the homogeneous case, we have proven that nonresetting in the bulk is the optimal strategy for certain target distributions, specifically those for which the region with positive $M(x_{T})$ (close to the walls) is more important, when weighted with the target distribution $P(x_{T})$, than the region with negative $M(x_{T})$ (close to the center). For the monoparametric family (12), this corresponds to $\beta<\beta_{c}$: Loosely speaking, the weight of targets close to the center is insufficient to destabilize the nonresetting strategy in the bulk. Indeed, this is physically sensible: Since the resetting boundaries already facilitate the search, extra resetting in the bulk is beneficial only if the weight of targets close to the center is large enough.
IV Optimization over heterogeneous resetting
Figure 3: Stability diagram of the nonresetting strategy for heterogeneous resetting. Positive (negative) values, represented as warm (cold) colors, thereof contribute to zero (nonzero) bulk resetting as the best search strategy. White color indicates saturated values. Top: Functional derivative $M(x_{T},x)$ defined in Eq. (IV). Bottom: Symmetrized functional derivative $M_{s}(x_{T},x)$ defined in Eq. (17) for the heterogeneous case. The contour line at which $M_{s}=0$ is marked (dashed line), as well as the value $x_{T}^{(0)}$ (black diamond) below which $M_{s}(x_{T},x)<0$, $\forall x$.
We now move to the general case with heterogeneous resetting $r(x)$. Similarly to the homogeneous case, whether or not bulk resetting is advantageous is controlled by the sign of the derivative of the average MFPT with respect to the resetting rate, evaluated at $r(x)\equiv 0$. For heterogeneous resetting, the key quantity is thus